OpenStudy (anonymous):

Solve 8 - 9(2x + 4) <= -2x +13. Write your solution using set builder notation. Please show all of your work.

OpenStudy (anonymous):

help please!!!!!!!!!!!!!!!!

OpenStudy (across):

First of all, don't hijack other people's threads. It is very rude, and discourages others from helping you. Secondly, I can't believe that you've asked over one hundred questions and have only contributed to the site once. Try and help someone when you can. Lastly, let\[8-9(2x+4)\leq-2x+13.\]Then\[8+18x-36\leq-2x+13,\\18x-28\leq-2x+13,\\20x\leq41,\\x\leq\frac{41}{20}.\]If you write this using set-builder notation, you get\[x=\left\{y\in\mathbb{R}:y\leq\frac{41}{20}\right\}.\]

OpenStudy (anonymous):

sorry. I didn't know I was being rude :(

OpenStudy (anonymous):

and thanks

OpenStudy (across):

I made a slight arithmetic mistake. Can you find it?

OpenStudy (anonymous):

I am NOT good at math at all. and I wouldn't want to confuse anyone by giving them the wrong answer. and again, I apologize!!

OpenStudy (across):

That's fine, and the correct answer should be:\[x=\left\{y\in\mathbb{R}:y\geq-\frac{41}{16}\right\}.\]

OpenStudy (anonymous):

Is the sign not to be turned the other way??

OpenStudy (anonymous):

ok. thought so!!